Status Report

NASA Space Station On-Orbit Status 10 Feb 2004

By SpaceRef Editor
February 10, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 10 Feb 2004

All ISS systems continue to function nominally, except as noted previously or below.

FE Alexander Kaleri worked several hours to install equipment for the Russian “Matryoshka” payload in the Service Module (SM). [After first setting up the individual hardware components, Kaleri installed the Matryoshka server (BSPN) on its holding frame behind a panel (#128) on the SM’s port side, then the power switching unit (BSK-5V) on its frame on the starboard side behind panel #121, tagging up with ground specialists as necessary. After the ground turned off power to the BITS2-12 onboard telemetry (TM) measurement system and its VD-SU control mode, Kaleri mated the TM connectors of both units to the BITS-212. Power and VD-Su were turned on again, and Sasha installed the cables of the onboard equipment control system (SUBA). The work will be continued tomorrow. Matryoshka collects radiation measurements in the SM and DC-1 docking compartment for studies of on-orbit radiation and long-term dose accumulation, using six SPD dosimeters deployed throughout the RS as well as by in a spherical Matryoshka-R “phantom”, which simulates a human torso. Note: Matryoshka is the name for the traditional Russian set of nestling dolls.]

CDR/SO Michael Foale started up the DOUG (Dynamic Operational Ubiquitous Graphics) application, followed by the 2-hr. proficiency training on the SSRMS (Space Station Remote Manipulator System). During the OBT (on-board training), he performed a walk-off of the robotarm from the MBS PDGF (Mobile Base System/Power & Data Grapple Fixture) to the Lab PDGF on top of the “Destiny” module, leaving the SSRMS positioned to view the EVA-9 on 2/26. [The ops started with a single joint maneuver to the Lab PDGF pre-grapple position, with clearances between SSRMS & Lab Camera being monitored with the MBS Mast camera. The Lab Camera was available to monitor the clearance between the SSRMS and the UHF antenna. Throughout the day the Canadarm2 remains outside a 3-ft. envelope of the surrounding structure (except during grapple and release operations). The inchworm-like base change walk-off was similar to one performed by Mike on his last MSS Ops day (12/18/03). The final maneuver of the day was a single joint maneuver to the EVA viewing position.]

Alex Kaleri unpacked the hardware kits of the Russian BIO-5 Rasteniya-2 (“plants-2”) experiment, set it up for operation and activated it by turning on environmental control power (pumps, light and fan), priming the tensiometers and setting laptop mode to cultivation. [Rasteniya-2 researches growth and development of plants under spaceflight conditions. After hardware installation, Kaleri planted six seeds of acacia leaf peas between the wicks of the root module, made power connections and locked the tray. Regular daily maintenance of the experiment involves monitoring of seedling growth, humidity measurements, moistening of the substrate if necessary, and photo/video recording.]

The CDR terminated the maintenance charge on the two new U.S. EMU/spacesuit batteries delivered on Progress 13P (#2045 & #2046) in the Airlock’s BSA (battery stowage assembly), then configured the controlling SSC7 (station support computer #7) laptop and initiated total discharge of the batteries. SSC7 was later reset to nominal configuration. [EMU (extravehicular mobility unit) battery maintenance is performed every 50 days, and consists of fully charging and then discharging the batteries to prolong their useful life.]

In preparation for the Lab science window depressurization activity tomorrow (2/11), Mike Foale worked near the window to install and tighten its QD (quick disconnect) coupling on its bracket. A crew conference with the ground was required before actual tightening the QD to clarify techniques and risks. [Carefully prepared procedures were uplinked for the very slow QD tightening/throttling technique designed to avoid flash condensation of any humidity in Volume D between the two pressure panes. Mike today first practiced the throttling technique by using male and female NOAPs (nitrogen/oxygen pressure adapters), an ISA (internal sample adapter) and VAJs (vacuum access jumpers) to slowly vent cabin air into an evacuated equipment volume. Tomorrow’s “real” depressurization will use the FSS (fluid system servicer), ISA and VAJ with the Lab VRIV (vent & relief isolation valve) to slowly vent the inter-pane “Volume D” overboard. ]

CDR Foale completed his twelfth weekly filling-out of the Food Frequency Questionnaire (FFQ), which keeps a log of his nutritional intake over time on the medical equipment computer (MEC).

Sasha performed the daily routine maintenance of the SM’s SOZh life support system.

Mike worked on the MSG (Microgravity Science Glovebox), tending the PromISS-3 (Protein Crystal Growth Monitoring by Digital Holographic Microscope 3) experiment by removing videotape #6 and installing tape #7.

Station attitude continues in LVLH YVV (local vertical/local horizontal, with y-axis in velocity vector, i.e., flying “sidewise”), until 2/21. ISS will then maneuver back to XPOP.

MCC-H flight controllers have performed a checkout of several unused RPCs (remote power controllers) on orbit, with the purpose to uncover any potentially degraded RPCMs (remote power controller modules) as early as possible. [Approximately 30% of the unused RPCs currently on orbit were tested yesterday (by closing and opening each unit), and the balance will be tested later this week. All of the RPCs tested to date operated properly.]

Today’s CEO (crew earth observations) targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Toshka Lakes, Egypt (looking right to see the whole line of new lakes in Egypt’s far south. These are changing rapidly as depressions in the desert fill with water), Saharan dust (the longest and most vigorous Saharan dust event this season continues to transport dust thousands of km westward from source points between the Atlantic coast and central Chad. On this pass over great bend of the Niger River, the crew was asked to look left and right for the northern margin of the dust mass [suggested by haze/visibility reports from local stations]. Timbuktu was left of track on the north side of the Niger), Ancient river courses, SW Libya (winding river courses formed under a vast ice sheet ~450 million years ago when the Sahara lay athwart the South Pole. The courses are known from a few localities. Remotely sensed imagery will assist in 1) discovering new sub-ice channels and 2) reconstructing the network of channels. Looking just right of nadir in the ancient [dark] rock masses), Salt Lake, Utah (lake levels of this long-term monitoring site appear at nadir and a right of track), Baker Island, Pacific (detailed mapping images of this remote, mile-long island were requested), and Howland Island, Pacific (looking left of track for this equatorial 1.5 mile long atoll entirely surrounded by a coral reef).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 1:07pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On. Vozdukh CO2 scrubber is On (in Automatic Mode). U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode). TCCS (trace contaminant control subsystem) is operating. SM Gas Analyzer has been calibrated and is used for ppO2 and ppCO2 monitoring. MCA (major constituents analyzer) is off (in Life Extending Mode). BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode. RS air conditioner SKV-1 is On; SKV-2 is Off (repair incomplete).
  • SM Working Compartment: Pressure (mmHg) — 742; temperature (deg C) — 25.5; ppO2 (mmHg) — 157.2; ppCO2 (mmHg) — 2.5;
  • SM Transfer Compartment: Pressure (mmHg) — 761; temperature (deg C) — 20.4.
  • FGB Cabin: Pressure (mmHg) — 756; temperature (deg C) — 23.0.
  • Node: Pressure (mmHg) — 747.71; temperature (deg C) — 23.4 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab: Pressure (mmHg) — 750.17; temperature (deg C) — 25.1; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 750.27; temperature (deg C) — 25.1; shell heater temp (deg C) — 24.5, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1: Shell heater temp (deg C) — 23.0
  • PMA-2: Shell heater temp (deg C) 11.0

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational. BGA (beta gimbal assembly) 2B and 4B both in Directed Position (2B: 235 deg; 4B: 125 deg); non-suntracking, “night glider”/”sun slicer” drag reduction mode.
  • SM batteries: Battery #8, formerly known as #7, is failed (to be replaced); battery #6 is in “Cycle” mode; all other batteries (6) are in “Partial Charge” mode.
  • FGB batteries: Battery #3 is off (capacity restoration mode, ROM); all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22).
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational (string #3 dropped out 10/22).

Propulsion System:

  • Total propellant load available: 4115 kg (9072 lb) as of 2/5/04 [SM(755) + FGB(2701) + Progress M(0) + Progress M-1(659)]. (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, y-axis in velocity vector [yaw: -90 deg, pitch: -9 deg, roll: 1.7 deg]), with CMG/TA (thruster assist) Momentum Management.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings.
  • MT: latched and mated at WS4.
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit (as of this morning, 8:04am EST [= epoch]):

  • Mean altitude — 367.6 km
  • Apogee — 372.7 km
  • Perigee — 362.6 km
  • Period — 91.9 min.
  • Inclination (to Equator) — 51.63 deg
  • Eccentricity — 0.0007429
  • Orbits per 24-hr. day — 15.67
  • Mean altitude loss last 24 hours — 80 m
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 29813

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see

SpaceRef staff editor.